OpenStudy (anonymous):

What is the magnitude and direction of if G(2, -2) and H(7, 6)? magnitude: 9.43 units; direction 32.01° magnitude: 6.4 units; direction: 57.99° magnitude: 9.43 units: direction: 57.99° magnitude: 6.4 units: direction: 32.01°

OpenStudy (anonymous):

magnitude = \[\sqrt{(7-2)^2+ (6-(-2))^2}=9.43\] direction = \[\tan^{-1} (6-(-2)/(7-2))= 57.99\]

OpenStudy (anonymous):

so C is correct

OpenStudy (anonymous):

ya absolutely

OpenStudy (anonymous):

thanks